# SynCoGen: Synthesizable 3D Molecule Generation via Joint Reaction and Coordinate Modeling

### v1

@@ Abstract
Ensuring synthesizability in generative small molecule design remains a major challenge. While recent developments in synthesizable molecule generation have demonstrated promising results, these efforts have been largely confined to 2D molecular graph representations, limiting the ability to perform geometry-based conditional generation. In this work, we present SynCoGen (Synthesizable Co-Generation), a single framework that combines simultaneous masked graph diffusion and flow matching for synthesizable 3D molecule generation. SynCoGen samples from the joint distribution of molecular building blocks, chemical reactions, and atomic coordinates. To train the model, we curated SynSpace, a dataset containing over 600K synthesis-aware building block graphs and 3.3M conformers. SynCoGen achieves state-of-the-art performance in unconditional small molecule graph and conformer generation, and the model delivers competitive performance in zero-shot molecular linker design for protein ligand generation in drug discovery. Overall, this multimodal formulation represents a foundation for future applications enabled by non-autoregressive molecular generation, including analog expansion, lead optimization, and direct structure conditioning.
@@ 1 Introduction
Generative models significantly enhance the efficiency of chemical space exploration in drug discovery by directly sampling molecules with desired properties. However, a key bottleneck in their practical deployment is low synthetic accessibility, i.e. generated molecules are often difficult or impossible to produce in the laboratory (Gao and Coley, 2020). To address this limitation, recent work has turned to template-based methods that emulate the chemical synthesis process by constructing synthesis trees that link molecular building blocks through known reaction templates (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025; Gao et al., 2024; Jocys et al., 2024; Swanson et al., 2024). These representations, while useful for downstream experimental validation, do not describe the underlying 3D geometry and thus cannot capitalize on the conformational information that is often crucial for diverse chemical and biological properties.
@@
This work introduces SynCoGen (Synthesizable Co-Generation), a generative modeling framework aiming to bridge the gap between 3D molecular generation and practical synthetic accessibility (Figure 1). Our main contributions are as follows:
@@
  • Molecular Dataset: We curate a new dataset SynSpace, comprising 622,766 synthesizable molecules represented as building block reaction graphs, along with 3,360,908 associated low-energy conformations. Compared to synthon-based datasets, this dataset enables models to generate more readily synthesizable molecules and directly suggest streamlined synthetic routes.
  • Empirical Validation: We demonstrate that SynCoGen achieves state-of-the-art performance in 3D molecule generation, producing physically realistic conformers while explicitly tracing reaction steps. Moreover, SynCoGen performs zero-shot 3D conditional molecular generation tasks such as linker design, highlighting its applicability for drug discovery.
Figure 1: SynCoGen is a simultaneous masked graph diffusion and flow matching model that generates synthesizable molecules in 3D coordinate space. Each node corresponds to a building block, and edges encode chemical reactions. Note that nodes are not necessarily linear and that the leaving groups are not displayed.
@@ 2 Background and Related Work
[…]
@@
The optimal reverse process $p_{\theta}(z_{s}\mid z_{t})$ takes the same form but with $x_{\theta}(z_{t},t)$ in place of the true $\mathbf{x}$. We adopt the zero-masking and carry-over unmasking modifications of Sahoo et al. (2024).
@@
Beyond directly optimizing synthesizability scores (Liu et al., 2022; Guo and Schwaller, 2025) – which are often unreliable – the predominant approach to ensuring synthetic accessibility involves modifying generative models to incorporate reaction templates. Early methods explored autoencoders (Bradshaw et al., 2019, 2020), genetic algorithms (Gao et al., 2022), and reinforcement learning (Gottipati et al., 2020; Horwood and Noutahi, 2020). More recently, GFlowNet-based (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025) and transformer-based (Gao et al., 2024; Jocys et al., 2024) methods have gained prominence. Such generative models have already shown practical utility in biological discovery tasks (Swanson et al., 2024). However, most methods only generate molecular graphs and do not produce 3D structures.
@@ 3 Dataset
Alternatively, many synthesis-aware generators employ external reaction simulators, such as RDKit, to couple building blocks iteratively. While convenient, such black-box steps offer no fine-grained control when a reagent has multiple reaction centers, distinct atoms or atom sets that can each serve as the specific site of bond formation or cleavage in a coupling reaction. They also do not define atom mappings between reactants and products, making it impossible to trace product atoms back to their parent building blocks, which complicates edge assignment in building block graph generation.
To overcome these limitations, we curate a new dataset SynSpace (Figure 2) comprising building block-level reaction graphs pairs with corresponding atom- and block-level graphs. We then calculate multiple 3D conformations for each graph using semi-empirical methods (Bannwarth et al., 2019). SynSpace is available at https://huggingface.co/datasets/DreiSSB/SynSpace.
Figure 2: Graphical overview of SynSpace creation process. Highly synthesizable molecules are procedurally constructed by sampling synthesis pathways from a predefined set of building blocks and reactions. Starting from an initial building block, the procedure selects a reaction center, a compatible reaction, and a suitable reactant. This process is iteratively repeated for a fixed number of reaction steps. After the final structure is assembled, multiple low-energy 3D conformations are generated.
@@ 3.1 SynSpace: Graph Generation
We begin by constructing a vocabulary of 93 commercially available, low-cost building blocks and 19 high-yield reaction templates. This vocabulary is adapted from the collection proposed by Koziarski et al. (2024), retaining reactions that (1) ensure all product atoms originate from the two input reagents, and (2) involve at most one leaving group per reagent. We emphasize that these are not only feasible chemistries, but also simple and reliable chemical reactions with readily available building blocks that can enable rapid multi-step synthesis.
We procedurally generate SynSpace from this vocabulary by iteratively coupling building block graphs at their respective reaction centers with compatible chemical reaction templates, described in detail in Section A.2. We obtained 622,766 building block reaction graphs, each constructed from 2 to 4 sequential reactions. For each resulting molecule, we generate multiple low-energy conformations and retain their atomic coordinates, with a total of 3,360,908 conformations.
@@
Many commercially available building blocks contain multiple reaction centers, each compatible with a different set of corresponding reaction centers on other building blocks. In this way, a building block-level reaction graph $G_{b}=(X,E)$ is not fully specified when edges are parametrized by the reaction alone. To achieve an injective correspondence, we therefore label edges from node $i$ to $j>i$ by the triple
$e_{ij}=(r,v_{i},v_{j}),$ (3)
where $r$ is the coupling reaction and $(v_{i},v_{j})$ are the participating reaction centers on the source and destination blocks, respectively. Strictly speaking, stereoisomers that form during reactions collapse to the same $(X,E)$ representation, but this granularity suffices for the scope of the current work.
@@ 3.2 SynSpace: Conformation Generation
For each molecular graph, 50 initial conformers were generated using the ETKDG (Riniker and Landrum, 2015) algorithm (RDKit implementation). These structures were energy-minimized using the MMFF94 force field, and all conformers within 10 kcal/mol of the global minimum were retained. The resulting geometries were then re-optimized with the semi-empirical GFN2-xTB (Bannwarth et al., 2019) method, after which the same 10 kcal/mol energy threshold was applied. At every stage, redundant structures were removed by geometry-based clustering ($\text{RMSD}<1.5\AA$). This workflow yields, on average, 5.4 distinct conformers per graph. Relative to exhaustive approaches such as CREST (Pracht et al., 2024), the workflow is several orders of magnitude faster; despite occasionally omitting some conformations, the retained structures are diverse and reproduce the bond-length, bond-angle, and dihedral-angle distributions observed in CREST-derived datasets (see Section 5.1).
@@ 4 Methods
Let $\mathcal{B}$ be the building-block vocabulary and $\mathcal{R}$ the set of reaction templates, with cardinalities $B:=|\mathcal{B}|$ and $R:=|\mathcal{R}|$. We write $N$ for the maximum number of building blocks that any molecule in the training set can contain, and $M$ for the maximum number of atoms in a single building block. For each block $b\in\mathcal{B}$ we denote its set of reaction-center atoms by $\mathcal{V}(b)$; the global maximum of these counts is $V_{\max}:=\max_{b\in\mathcal{B}}\lvert\mathcal{V}(b)\rvert$.
Hence, tensor shapes contain factors such as $B+1$ (to accommodate the masked token $\pi_{X}$ in $X$), $R\,V_{\max}^{2}+2$ (to accommodate the no-edge and masked tokens $\lambda_{E}$ and $\pi_{E}$), together with the bounds $N$ and $M$ introduced above. For any coordinate tensor $C$ and binary mask $S$ we define the mask-weighted centroid and its centered version by $\bar{C}_{S}:=\frac{\sum S\odot C}{\sum S},\qquad\tilde{C}:=C-\bar{C}_{S}.$
@@
SynCoGen generates building block-level reaction graphs and coordinates. Each molecule is represented by a triple $(X,E,C)$ where $X\in\{0,1\}^{N\times|\mathcal{B}|+1}$ encodes the sequence of building-block identities, $E\in\{0,1\}^{N\times N\times|\mathcal{R}|V_{\max}^{2}+2}$ labels the coupling reaction (and centers) between every building block pair, and $C\in\mathbb{R}^{N\times M\times 3}$ stores all atomic coordinates. We detail the parameterization of graphs $(X,E)$ in Section B.1. Training combines two diffusion schemes: 1) a discrete absorbing process on $(X,E)$ using the categorical forward kernel of Sahoo et al. (2024), and 2) a continuous, visibility-aware process on $C$ whose endpoints are (i) a rototranslationally-aligned isotropic Gaussian and (ii) a re-centered ground truth, considering all "visible" atoms in the prior (see Section 4.2). The code is available at https://github.com/andreirekesh/SynCoGen.
@@ 4.1 Model Architecture
At each timestep $t$, SynCoGen predicts building block logits $L_{t}^{X},L_{t}^{E}$ and a shifted coordinate estimate $\hat{\tilde{C}}^{\,t}_{0}$. The total loss is the weighted sum of the cross-entropy term $\mathcal{L}_{\mathrm{graph}}$ on $(X,E)$, the masked coordinate MSE term $\mathcal{L}_{\mathrm{MSE}}$, and the short-range pairwise distance term $\mathcal{L}_{\mathrm{pair}}$ (see Section B.3 for the training algorithm and Section B.11 for loss details). We implement a modified version of SemlaFlow (Irwin et al., 2025), a $SE(3)$ equivariant architecture originally designed for all-atom molecular design, as the principal backbone to generate both coordinates and graphs. As SemlaFlow handles atom and atom-level bond representations as both inputs and outputs, we define additional building-block-to-atom featurization protocols (Section B.2) and atom-to-building-block output layers (Section B.5).

### v2

@@ Abstract
Synthesizability remains a critical bottleneck in generative molecular design. While recent advances have addressed synthesizability in 2D graphs, extending these constraints to 3D for geometry-based conditional generation remains largely unexplored. In this work, we present SynCoGen (Synthesizable Co-Generation), a single framework that combines simultaneous masked graph diffusion and flow matching for synthesizable 3D molecule generation. SynCoGen samples from the joint distribution of molecular building blocks, chemical reactions, and atomic coordinates. To train the model, we curated SynSpace, a dataset family containing over 1.2M synthesis‑aware building block graphs and 7.5M conformers. SynCoGen achieves state-of-the-art performance in unconditional small molecule graph and conformer co-generation. For protein ligand generation in drug discovery, the amortized model delivers superior performance in both molecular linker design and pharmacophore-conditioned generation across diverse targets without relying on any scoring functions. Overall, this multimodal non-autoregressive formulation represents a foundation for a range of molecular design applications, including analog expansion, lead optimization, and direct de novo design.
@@ 1 Introduction
Generative models significantly enhance the efficiency of chemical space exploration in drug discovery by directly sampling molecules with desired properties. However, a key bottleneck in their practical deployment is low synthetic accessibility; that is, generated molecules are often difficult or impossible to produce in the laboratory (Gao and Coley, 2020). To address this limitation, recent work has turned to template-based methods that emulate the chemical synthesis process by constructing synthesis trees that link molecular building blocks through known reaction templates (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025; Gao et al., 2024; Jocys et al., 2024; Swanson et al., 2024). These representations, while useful for downstream experimental validation, do not describe the underlying 3D geometry and thus cannot capitalize on the conformational information that is often crucial for diverse chemical and biological properties.
@@
This work introduces SynCoGen (Synthesizable Co-Generation), a generative modeling framework aiming to bridge the gap between 3D molecular generation and practical synthetic accessibility (Figure˜1). Our main contributions are as follows:
@@
  • Molecular Dataset: We curate a new dataset family SynSpace, comprising 1.2M synthesizable molecules represented as building block reaction graphs, along with 7.5M associated low-energy conformations. Compared to synthon-based datasets, SynSpace enables models to generate more readily synthesizable molecules and directly suggest streamlined synthetic routes.
  • Empirical Validation: We demonstrate that SynCoGen achieves state-of-the-art performance in 3D molecule generation, producing physically realistic conformers while explicitly tracing reaction steps. Ablations show our modelling choices are crucial for the performance. Importantly, SynCoGen performs 3D conditional molecular generation tasks including linker design and pharmacophore-conditioned generation, highlighting its applicability for drug discovery.
Figure 1: SynCoGen is a simultaneous masked graph diffusion and flow matching model that generates synthesizable molecules in 3D coordinate space. Each node corresponds to a building block, and edges encode chemical reactions. Note that graphs are not necessarily path graphs, the leaving groups are not displayed, and there is no order to which nodes and edges are denoised.
@@ 2 Background and Related Work
Let $\mathbf{x}\sim\rho_{\text{data}}$ be a one-hot encoding over $K$ categories. Discrete diffusion models (Austin et al., 2021; Sahoo et al., 2024; Shi et al., 2024) map the complex data distribution $\rho_{\text{data}}$ to a simpler distribution via a Markov process, with absorbing (or masked) diffusion being the most common. In the masked diffusion framework, the forward interpolation process $(\rho_{t})_{t\in[0,1]}$ with the associated noise schedule $(\alpha_{t})_{t\in[0,1]}$ results in marginals $q(\mathbf{z}_{t}|\mathbf{x})=\text{Cat}(\mathbf{z}_{t};\alpha_{t}\mathbf{x}+(1-\alpha_{t})\mathbf{m})$, where $\mathbf{z}_{t}$ and $\mathbf{m}$ denote intermediate latent variables and the one-hot encoding for the special $\text{MASK}]$ token, respectively. The posterior can be derived as:
@@
The optimal reverse process $p_{\theta}(z_{s}\mid z_{t})$ takes the same form but with $x_{\theta}(z_{t},t)$ in place of the true $\mathbf{x}$. We adopt the zero‐masking and carry‐over unmasking modifications of Sahoo et al. (2024).
@@
Beyond directly optimizing synthesizability scores (Liu et al., 2022; Guo and Schwaller, 2025) – which are often unreliable – the predominant approach to ensuring synthetic accessibility in generative models is to incorporate reaction templates. Early methods explored autoencoders (Bradshaw et al., 2019; 2020), genetic algorithms (Gao et al., 2022), and reinforcement learning (Gottipati et al., 2020; Horwood and Noutahi, 2020). Recently, GFlowNet-based (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025) and transformer-based (Gao et al., 2024; Jocys et al., 2024) methods have gained prominence. Such generative models have already shown practical utility in biological discovery tasks (Swanson et al., 2024). However, most methods only generate molecular graphs and do not produce 3D structures. The recent CGFlow Shen et al. (2025) performs 3D generation via a GFlowNet policy augmented with flow matching; however, CGFlow optimizes a reward and typically requires a full training for each target pocket.
@@ 3 Dataset
Alternatively, many synthesis‑aware generators employ external reaction simulators, such as RDKit, to couple building blocks iteratively. While convenient, such black‑box steps offer no fine‑grained control when a reagent has multiple reaction centers, distinct atoms or atom sets that can each serve as the site of bond formation or cleavage in a reaction. They also do not define atom mappings between reactants and products, making it impossible to trace product atoms back to their parent building blocks, which complicates edge assignment in building block graph generation. To overcome these limitations, we curate a new family of datasets, SynSpace (Figure˜2), comprising building block-level reaction graphs pairs with corresponding atom- and block-level graphs. We then calculate multiple 3D conformations for each graph using semi-empirical methods (Bannwarth et al., 2019).
Figure 2: Overview of SynSpace creation process. Highly synthesizable molecules are procedurally constructed by iteratively sampling synthesis pathways from a set of building blocks and reactions. Starting from an initial block, the procedure selects a reaction center, a compatible reaction, and a suitable reactant. After the final structure is assembled, multiple low-energy 3D conformations are generated. We provide two SynSpace datasets from two vocabularies, a practically focused core set and an extended variant; each dataset contains 600k graphs with 3-4M conformers.
@@ 3.1 SynSpace: Graph Generation
We first construct two curated vocabularies adapted from the collection proposed by Koziarski et al. (2024). The first vocabulary pairs 93 low-cost, commercially available building blocks with 19 high-yield reaction templates, defining a virtual synthesis space of over a billion molecules. The second vocabulary is a superset with 378 building blocks with 26 reactions, expanding the synthesis space to over a trillion molecules. All building blocks were selected because they are known to undergo the chosen reactions, acknowledging that the presence of a nominally compatible functional group alone does not guarantee participation in the corresponding transformation. We utilize reactions that (1) ensure all product atoms originate from the two input reagents, and (2) involve at most one leaving group per reagent. We emphasize that these constraints yield simple, robust chemistries that are routinely executed and support rapid multi-step synthesis from inexpensive, in-stock reagents.
We procedurally generate SynSpace from the smaller vocabulary, or SynSpace-L from the larger superset, by iteratively coupling building block graphs at their reaction centers with compatible reaction templates (Section˜A.2). For SynSpace, we obtained 622,766 building block reaction graphs, each constructed from 2 to 4 sequential reactions. For each molecule, we generate multiple 3D conformations (Section˜3.2), yielding 3,360,908 conformers. Similarly, SynSpace-L contains 600,000 graphs and 4,223,367 conformations. Unless otherwise noted, all models are trained on SynSpace, which emphasizes practicality as its fewer building blocks are more readily stocked, whereas SynSpace-L is reserved for when a larger, more exploratory search space is required.
SynSpace contains diverse molecules that are drug-like (e.g., LogP $\sim 2.5$; broad range of topological polar surface areas; large fraction of sp3 carbons). Importantly, compared to Geom-Drugs (Axelrod and Gomez-Bombarelli, 2022), SynSpace contains substantially more unique Murcko scaffolds, indicating breadth despite the building block space. With a larger accessible space, SynSpace-L preserves similar physicochemical profiles and scaffold diversity. See Section˜A.3 for details.
@@
Many commercial building blocks contain multiple reaction centers, each compatible with a different set of corresponding reaction centers on other blocks. Thus, a building block-level reaction graph $G_{b}=(X,E)$ is not fully specified when edges are parametrized by the reaction alone. To achieve an injective correspondence, we label edges from node $i$ to $j>i$ by the triple $e_{ij}=(r,v_{i},v_{j})$, where $r$ is the coupling reaction and $(v_{i},v_{j})$ are the participating reaction centers on the source and destination blocks, respectively. Stereoisomers that form during reactions collapse to the same $(X,E)$ representation, but this granularity suffices for our current scope.
@@ 3.2 SynSpace: Conformation Generation
For each molecular graph, 50 initial conformers were generated using the ETKDG (Riniker and Landrum, 2015) algorithm (RDKit implementation). These structures were energy-minimized using the MMFF94 force field, and all conformers within 10 kcal/mol of the global minimum were retained. The resulting geometries were then re-optimized with the semi-empirical GFN2-xTB (Bannwarth et al., 2019) method, after which the same 10 kcal/mol energy threshold was applied. At every stage, redundant structures were removed by geometry-based clustering ($\text{RMSD}<1.5\text{\AA }$). This workflow yields, on average, 5.4 distinct conformers per graph. Relative to exhaustive approaches such as CREST (Pracht et al., 2024), the workflow is several orders of magnitude faster; despite occasionally omitting some conformations, the retained structures are diverse and reproduce the bond-length, bond-angle, and dihedral-angle distributions observed in CREST-derived datasets (see Section˜5.1).
3.3 SynSpace: Pharmacophore Generation
For each conformer of a molecule in SynSpace and SynSpace-L, we generate a pharmacophore profile consisting of one-hot pharmacophore types $X_{\text{pharm}}\in\{0,1\}^{N_{\text{pharm}}\times N_{\text{types}}}$ and positions $C_{\text{pharm}}\in\mathbb{R}^{N_{\text{pharm}}\times 3}$ using ShePhERD Adams et al. (2025). Here, $N_{\text{pharm}}$ and $N_{\text{types}}$ correspond to the number of pharmacophore features and the number of pharmacophore types, respectively.
@@ 4 Methods
Let $\mathcal{B}$ be the building‑block vocabulary and $\mathcal{R}$ the set of reaction templates, with cardinalities $B:=|\mathcal{B}|$ and $R:=|\mathcal{R}|$. Let $N$ denote the maximum number of building blocks per molecule and $M$ the maximum number of atoms per building block. For each block $b\in\mathcal{B}$ we denote its set of reaction‑center atoms by $\mathcal{V}(b)$; the global maximum of these counts is $V_{\max}:=\max_{b\in\mathcal{B}}\lvert\mathcal{V}(b)\rvert$. Hence, tensor shapes contain factors such as $B+1$ (to accommodate the masked token $\pi_{X}$ in $X$), $R\,V_{\max}^{2}+2$ (to accommodate the no-edge and masked tokens $\lambda_{E}$ and $\pi_{E}$), together with the bounds $N$ and $M$ introduced above.
@@
SynCoGen generates building block-level reaction graphs and coordinates. Each molecule is represented by a triple $(X,E,C)$ where $X\in\{0,1\}^{N\times|\mathcal{B}|+1}$ encodes the sequence of building‑block identities, $E\in\{0,1\}^{N\times N\times|\mathcal{R}|V_{\max}^{2}+2}$ labels the coupling reaction (and centers) between every building block pair, and $C\in\mathbb{R}^{N\times M\times 3}$ stores all atomic coordinates. We detail the parameterization of graphs $(X,E)$ in Section˜B.4. Training combines two diffusion schemes: 1) a discrete absorbing process on $(X,E)$ using the categorical forward kernel of Sahoo et al. (2024), and 2) a continuous, visibility‑aware process on $C$ whose endpoints are (i) a rototranslationally‑aligned isotropic Gaussian and (ii) a re-centered ground truth, considering all "visible" atoms in the prior (see Section˜4.2). For an intuitive schematic and description of the training procedure, see Section˜B.1.
@@ 4.1 Model Architecture
We adapt a $SE(3)$-equivariant architecture originally designed for all-atom molecular design (SemlaFlow (Irwin et al., 2025)) as the principal backbone to generate both coordinates and graphs. At each timestep $t$, SynCoGen predicts building block logits $L_{t}^{X},L_{t}^{E}$ and a shifted coordinate estimate $\hat{\tilde{C}}^{\,t}_{0}$. The loss is the weighted sum of the cross-entropy term $\mathcal{L}_{\mathrm{graph}}$ on $(X,E)$, the masked coordinate MSE term $\mathcal{L}_{\mathrm{MSE}}$, and the short-range pairwise distance term $\mathcal{L}_{\mathrm{pair}}$ (see Sections˜B.7 and B.16 for details). We define additional building-block-to-atom featurization in Section˜B.5 and atom-to-building-block output layers in Section˜B.10.
Pharmacophore Conditioning Backbone.
To accommodate pharmacophores as conditioning information, we design a modified backbone to represent each as an "atom" with no weight during centering operations. After atom featurization, pharmacophore types are fed through a separate featurization head and concatenated to invariant atom type features, i.e. $X_{\text{model}}=[\text{MLP}_{\text{atom}}(X_{\text{atom}}),\,\text{MLP}_{\text{pharm}}(X_{\text{pharm}})]\in\mathbb{R}^{(N+N_{\text{pharm}})\times d_{x}}$. Pharmacophore coordinates are concatenated directly to atomic coordinates, $C_{\text{model}}=[C,C_{\text{pharm}}]\in\mathbb{R}^{(N+N_{\text{pharm}})\times 3}$, and therefore undergo identical data augmentation beforehand (including that induced by data pairing, see Section 4.2). $C_{\text{model}}$ and $X_{\text{model}}$ are then passed to the equivariant-invariant dynamics module. Prior to final output layers, expanded atom-level hidden-layer outputs are truncated to the total number of atoms $NM$.